The value derived from exercising an option immediately is the exercise value. No rational investor would exercise an option that is out-of-the-money, so the minimum exercise value is zero.

The following table provides information regarding options on ABC Corp. stock. Because the stock’s price is volatile, investors trade options to either hedge their positions or speculate on price movements. Investors can either buy options or “issue” new options, which is called writing options.

Consider the case of ABC Corporation:

The following table presents the data on ABC Corp.’s call options at different stock prices. Based on your understanding of exercise value and option prices, complete the table with a strike price of $28.00:

Stock Price ($)

Strike Price ($)

Exercise Value ($)

Market Price of Option ($)

Time Value ($)

8.00 28.00 0.00 1.56   
16.00 28.00    2.10 2.10
20.00 28.00    2.40 2.40
22.00 28.00 0.00 2.60   
24.00 28.00    4.00 4.00

After two weeks, the stock price of ABC Corp. increases to $24.96. Suppose you purchased the shares for $16.00 and then sell the shares at $24.96; your rate of return will be   on that transaction.

After your analysis, you decided to purchase the option for $2.10, which gives you the right to buy ABC Corp.’s stock at $28.00. If you exercise the option by purchasing the stock at the strike price, you could immediately sell the share of the stock at its market price of $24.96. This will result in a payoff, and the rate of return on your option will be   .

Answer the following questions, based on your understanding of options, option price, and exercise value.

True or False: Your friend thinks that the stock price of KnowItAll Corp. will decline. He decides to write and sell an option without buying the stock. He says that he will buy the stock when his option buyer exercises the option. This is an example of a naked option.
This statement is   .
True or False: Corporations on whose stocks options are traded in the option market are directly involved in the process, and it helps them raise additional funds.
This statement is   .
True or False: Options can be created and traded without an underlying asset.
This statement is   .
Answer #1

For Call Option , Intrinsic value = max(Price of Underlying Asset - Strike Price, 0)

Time value = Premium - Intrinsic Value

Stock Price ($) Strike Price ($) Exercise Value ($) Market Price of Option ($) Time Value ($)
8 28 0 MAX(8-28,0) 1.56 1.56 1.56-0
16 28 0 MAX(16-28,0) 2.1 2.1 2.1-0
20 28 0 MAX(20-28,0) 2.4 2.4 2.4-0
22 28 0 MAX(22-28,0) 2.6 2.6 2.6-0
24 28 0 MAX(24-28,0) 4 4 4-0

Answer 2)

In case of Spot market :  

purchased the shares for $16.00 and then sell the shares at $24.96

return on transaction : 24.96 -16.00/ 16.00 = 56 % in 2 weeks

annualized rate of return by using the formula of r= 56%*26 = 1456%

in case of call option exercise ,Payoff = Max( 24.96-28,0)=0

return on transaction : 0-2.10/2.10 = -100 % in 2 weeks

Answer 3)

True or False: Your friend thinks that the stock price of KnowItAll Corp. will decline. He decides to write and sell an option without buying the stock. He says that he will buy the stock when his option buyer exercises the option. This is an example of a naked option.
This statement is   .TRUE
True or False: Corporations on whose stocks options are traded in the option market are directly involved in the process, and it helps them raise additional funds.
This statement is FALSE
True or False: Options can be created and traded without an underlying asset.
This statement is   .FALSE